# Electrons herald non-classical light

### v1

@@ Abstract
To test the non-classicality of the radiation process, i.e., whether it is incompatible with classical field theory, we invoke the Cauchy-Schwarz inequality (CSI) [65] $\gamma=\frac{g_{E\textrm{A}\cup\textrm{B}}^{(2)}(\tau)}{g_{E}^{(2)}g^{(2)}(0)}\leq 1$, assuming a classical radiation process. The inequality (for derivations, see SI) bounds the temporal cross-correlation between the detected electron energy and photon number $g_{E,\textrm{A$\cup$B}}^{(2)}(\tau)$ to the variation of the electron energy loss $g_{E}^{(2)}=\frac{\braket{\Delta E_{el}^{2}}}{\braket{\Delta E_{el}}^{2}}$ and the temporal intensity correlation of photons in a Hanbury-Brown and Twiss setup $g^{(2)}(\tau)$, normalized to $\tau\gg 0$. For electron-photon coincidence, we find a pronounced violation of the CSI ($\gamma>1$, cf. Fig. 4c), which proves the non-classical nature of the process [66, 67, 68].
@@
This photon statistics of this radiation is further examined by the temporal intensity correlation $g^{(2)}$ of detectors placed in a Hanbury-Brown Twiss (HBT) setup (see Methods). While classical light shows $g^{(2)}(0)\geq g^{(2)}(\tau)$, this is not the case for non-classical light such as single photons or higher photon number states. The unheralded photon-photon intensity correlation (taken with higher time-resolution, higher-efficiency detectors and a time tagger, see Methods) shows a bunching peak at $\tau=0$ ps [69], inversely proportional to the electron current (cf. Fig. 4d). Following a current-dependent bunching height $g^{(2)}(0)\sim 1/(I\tau_{\mathrm{bin}})$, where $I$ is the electron flux rate and $\tau_{\mathrm{bin}}$ is the temporal data bin width, this behavior is expected when single electrons can create multiple photons, either in a direct process, as demonstrated here (see SI for further calculations), or as a result of cascaded excitation of multiple two-level systems leading to incoherent emission [70, 24, 71, 72].
@@
The statistics of the electron-heralded photon state are analyzed using the heralded intensity correlation (see Methods)
@@
relating the occurrence of electron-photon-photon coincidences $N_{\textrm{e,A,B}}$ (cf. Fig. 4a,b) to their single-photon equivalents $N_{\textrm{e,A/B}}$ and the number of heralding electrons $N_{\textrm{e}}$. For any energy window $E_{\textrm{el}}$, $g^{(2)}_{H}(\tau_{\textrm{A}},\tau_{\textrm{B}})$ is normalized to single-photon generation ($\tau_{A/B}\approx 0$, $\tau_{B/A}\neq 0$), thus uncovering true threefold coincidences ($\tau_{\textrm{A$\cap$B}}=0$).
@@
Figure 4e displays $g^{(2)}_{H}(\tau_{\textrm{A}},\tau_{\textrm{B}})$ for single-photon electron-energy losses (${m=1}$), as well as the so-called time-averaged heralded intensity correlation $g_{H}^{(2)}(\tau)$ reduced to the relative photon delay (inset; see SI for calculation) [73, 74]. The background level and the diagonal feature ($\tau_{\textrm{A$\cap$B}}\neq 0$) originate from accidental coincidences (cf. Fig. 4a). For electron-photon-photon coincidences ($\tau_{\textrm{A$\cap$B}}=0$), $g^{(2)}_{H}$ drops sharply to near-zero values, as also visible in the inset. Hence, for $m=1$, two photons detected at similar times are generated by two different electrons. Usually, accidental and true heralded photon-photon coincidences cannot be distinguished in the heralded intensity correlation if the coherence time is lower than the respective detector jitters [74, 75] (see also SI). However, in our case, the electron-heralded photon-photon coincidences remain distinct due to the much better relative photon detection jitters (see SI). A discrete version of the intensity correlation function $g^{(2)}_{H}[q]$ substitutes the time delay $\tau$ with the number of heralding electrons $q$ between two heralded photons. Here, the detection of two photons from a single electron is compared with random coincidences of two photons heralded by sequential electrons (for details, see Methods). The resulting intensity correlations are shown in Fig. 4f. Without energy filtering (top), a small bunching peak is observed $g^{(2)}_{H}[0]=1.10$, similar to the unheralded intensity correlation function (cf. Fig. 4d), but with larger bins corresponding to the electron detector resolution. Filtering to a single-quantum energy exchange ($m=1$), however, leads to a strong anti-bunching behavior with $g^{(2)}_{H}(0)$ suppressed to 0.056. The pronounced anti-bunching dip observed for both the time-dependent and discrete intensity correlation functions confirms the high purity of the heralded single-photon state. Two-photon states should, under ideal conditions, also show prominent antibunching. At present, however, technical aspects, including optical losses, electron spectral mixing, and the statistics of the electron source limit the ultimate purity of the heralded state (see SI). Such limitations can be improved upon by time-gated electron sources and enhanced sample designs.
@@
Quantifying our scheme in terms of electron-heralded photon generation requires a separation between different electron scattering pathways. We observe inelastic electron scattering at the material, as well as the parametric photon generation into broadband spatial optical modes. While the electron spectrum shows a mean coupling constant of $g_{0}^{\textrm{EELS}}>0.3$, the scattering into the single detected optical mode features a slightly smaller coupling with $g_{0}>0.2$ (cf. Fig. 3). This is the strongest observed coupling of electrons to a detector-coupled photonic mode observed thus far, facilitated by phase-matched, spatially extended electron-light interaction [36, 34, 76, 77]. Reaching the strong-coupling regime by establishing unity coupling constant to a single optical mode [50, 78, 79, 80, 55] will give direct access to a wide class of new phenomena, such as electron-photon entanglement [81, 82, 83, 84], photon-induced electron-electron correlations [33, 85] and the generation of more complex non-Gaussian optical states [53].
@@
In conclusion, we introduce electron-heralded quantum states of light, identifying hallmark features such as antibunching photon statistics and multiphoton-electron correlations. The scheme will facilitate electron-heralded high-order Fock states, based on further technical optimizations and strong single-mode coupling. Combined with photon or electron state operations, the generation of a wide variety of complex quantum states comes into reach. Broadening the scope of free-electron quantum optics, our results experimentally establish electron beams as a tailored quantum optical resource.
@@ Methods
The photonic chip with a $\text{Si}_{3}\text{N}_{4}$ waveguide embedded in $\text{SiO}_{2}$ is fabricated via the Damascene process [39] onto a Si substrate. The straight part of the waveguide (width x height $2.2\mu\text{m}$ x $780$ nm) interacting with the electron beam is $80\leavevmode\nobreak\ \mu\text{m}$ in length and has no top cladding, allowing for the optical modes to reach into the vacuum and interact with the electron beam. An ITO (indium tin oxide) layer of $<20$ nm thickness (index of refraction $n=1.99$ at 1550 nm measured at ITO films deposited under similar conditions) on top of the waveguide reduces charging in the presence of an electron beam. The chip is mounted on a custom-made TEM holder [40] such that the waveguide sits in parallel to the electron beam. The optical outputs of the symmetrically-built chip are connected to two short pieces of ultra-high numerical aperture (UHNA-7) fiber, which are spliced to single-mode fibers (SMF-28). The single-mode fibers are sent through the hollowed-out TEM holder and a custom vacuum fiber feedthrough. As the photon generation is directional, the optical detection setup is spliced to the SMF-28 fiber, which is connected to the lower output port.
@@
The experiments are conducted at the Göttingen UTEM (JEOL JEM 2100F) with a Schottky field-emission electron source operated in a continuous (thermal-field) mode. The resulting electron beam shows an energy spread of around 0.6 eV at an electron energy of 100 keV. The current is reduced to around 1.4 pA by adjusting the emitter temperature and by using a condenser aperture of $40\leavevmode\nobreak\ \mu\text{m}$ diameter to not oversaturate the electron detector. The experiments are conducted in low-magnification STEM (scanning transmission electron microscopy) mode, enabling the positioning and scanning of the focused electron beam (30 nm beam diameter, 1.1 mrad half convergence angle) in the sample plane. The scanning resolution is on the order of 30 nm. All measurements described in the main text are conducted with a stationary electron beam, which is checked and manually readjusted for an optimal combination of photon generation rate and electron transmission onto the detector every 10-30 min. Transmission losses on the electron side arise from charging effects along the waveguide surface, distorting and shifting the electron beam, which reduces incoupling into the 5 mm spectrometer entrance aperture. Post-measurement calculations show an electron transmission efficiency of around 60-70%. For the scanned data, the electron beam is scanned multiple times across the same region (width x height 200x500nm, 300 ms integration per pixel, pixel size $\sim 40$ nm). The scanned data is used to measure the position-dependent change of the coupling constant.
@@
The generated photons are sent through the waveguide and optical fibers to the outside of the TEM holder. The fiber end is spliced to a 50/50 beam splitter (Thorlabs TW1550R5A1), which distributes the photons to two cooled avalanche photodiodes (ID Quantique ID230). As the generated light is measured to be around $1300-1400\text{nm}$, we assume slightly higher losses and a deviation from the 50/50 coupling ratio at the beam splitter. The count rates indicate a splitting ratio of 52/48. The avalanche photodiodes are set to $>25\%$ detection efficiency and $20\mu\text{s}$ dead time, which leads to 250 and 300 photons/s intrinsic dark counts, respectively, for diode A and B. An evaluation of the transmission and detection efficiency shows an approximate efficiency of detecting a generated photon of 2% per diode. The optical spectrum is measured using a 10 nm flat-top bandpass filter (WL Photonics, tunability from 1200 to 1600nm), which is set between the TEM holder and the beam splitter, connected to the optical fibers via FC/APC connectors. The transmission band was moved incrementally between 1270 and 1570nm and the count rate on the detector was measured for every position (see SI). The spectrum shows a central wavelength of 1380 nm and a bandwidth of 100 nm (cf. Fig3e). The unheralded photon-photon intensity correlation is measured with a combination of superconducting nanowire single-photon detectors (Single Quantum) and a time tagger (Swabian Instruments, Time Tagger Ultra). The nanowire detectors show a detection efficiency $>80\%$ and a time resolution $<20$ ps and, therefore, allow for sharper temporal correlation when operated in combination with a Tagger unit with an RMS jitter of 9 ps.
@@
The electrons are separated in energy via a spectrometer (CEOS GmbH, CEFID) and detected on a hybrid pixel detector based on a Timepix ASID (Amsterdam Scientific Instruments, EM CheeTah T3). Every electron is categorized by its position on the chip (514x514 pixel, 0.03eV energy bin per pixel) and time-of-arrival (1.56 ns resolution). The stream of data consists of the electron data and the information from both time-to-digital converters at which the photon detector signals arrive. The photon detector signals are time-stamped with a 260 ps bin width and synchronized to the electron time stamps via the same global ASID clock. This allows for temporal correlation of the electron and photon arrival. Further post-acquisition steps are conducted to improve temporal and energy resolution on the electron side: First, a calibration of the individual pixel response times is done with regard to a mean detector response to enhance time resolution [36]. Second, single-particle clustering of the multiple excited pixels with a maximum cluster size of 10 gives access to individual electrons. Their localization follows the arithmetic mean position and the earliest arrival time inside a single cluster. A comparison between external current measurements and event counting on the detector shows an average of 3.4 excited pixels per electron. Last, time-dependent jitters of the electron beam position are reduced on a 10s interval by correcting the position of the zero-loss peak maximum.
@@
For the temporal correlation between electrons and photons, every detected electron is connected to the temporally closest photons on detector A and detector B, and the delay between the arrival times is calculated. The time offset due to propagation delays is measured via the peak position of the electron-photon A, electron-photon B, and photon A-photon B coincidence peaks. These offsets are subtracted for the depicted plots. The connection of electrons with two photons leads to a data set consisting of electron-photon-photon events with electron position on the detector, arrival times of all three particles, as well as time delays between them. For scanned data, the scan position is also included. The electron energy is determined from the position of the zero-loss peak center, which defines the position of 0 eV energy loss, and calibration images of the unscattered electron beam, which give access to the dispersion along the detector.
The electron-photon-photon events can then be analyzed as a data cube with the axes being the two time delays $\tau_{\textrm{A}}=t_{\textrm{A}}-t_{\textrm{el}}$ and $\tau_{\textrm{A}}=t_{\textrm{A}}-t_{\textrm{el}}$, as well as the electron energy loss with respect to to the uncorrelated zero-loss-peak (ZLP) position. A general fit function (see SI) can describe all features present in the cube. For easier access, individual cuts along or integrals over a specific axis are also analyzed and fitted independently with the fit function defined by the cube fit, as described in the main text.
@@
All data besides the uncorrelated spectrum (Fig. 2) is analysed filtered to a maximum electron-photon time delay of 100 ns, which strongly reduces computation time.
The generation of electron-photon-photon events, combined with large electron currents, can lead to multiple electrons being in coincidence with a single photon. Thus, if multiple electrons are temporarily close to a detected photon, the pair with the smallest time difference is considered a true coincidence, preventing multi-indexing of photons. These true coincidences are used for the analysis of the electron-heralded photon-photon correlation and the discrete intensity correlation.
@@
The statistics of photons and the proof for single photon states are often given via the intensity correlation measured in a Hanbury-Brown and Twiss setup: The photons are separated via a 50/50 beam splitter and sent on two photon detectors. Their signal is cross-correlated with each other and normalized to large time delays [63, 86]
@@
where $P_{\textrm{A,B}}(\tau)$ is the probability of detector A and B being excited at a temporal delay $\tau$ to each other, and $P_{\textrm{i}}$, i=A,B is the probability of detector excitation A or B. In the given experimental setup, the equation is adjusted for correlation counts $N$ instead of coincidence probabilities $P_{\textrm{i}}=N_{\textrm{i}}\tau_{\textrm{b}}/T_{\textrm{meas}}$ by adding the length of coincidence bins $\tau_{\textrm{b}}$ and the total measurement time $T_{\textrm{meas}}$.
As the electrons do not arrive uniformly, their statistical arrival distribution plays a role in the unheralded photon-photon intensity correlation. To remove the electron influence, an electron-heralded intensity correlation is employed. Heralded intensity correlation follows [73]
@@
in which the photon-photon cross-correlation $P_{\textrm{A,B,e}}(t_{\textrm{A}},t_{\textrm{B}},t_{\textrm{e}})$ is limited to a simultaneous detection of a heralding particle $e$ (see SI for more information). The function is again normalized to large photon-photon time delays, but under the condition that either photon A or photon B is in coincidence with a herald $e$. For electron-heralding intensity correlation the equation is rewritten in terms of correlation counts $N$ instead of coincidence probabilities, giving rise to Equation 3.
Due to the addition of a third particle, the heralded intensity correlation is a function of two time delays $\tau_{\textrm{A}}$ and $\tau_{\textrm{B}}$. A reduction to one single variable is done in time-averaged intensity correlation [73, 74]
$g^{(2)}(\tau_{\textrm{A}}=0,\tau=\tau_{\textrm{B}})=\frac{P_{\textrm{A,B,e}}(\tau)P_{\textrm{e}}}{P_{\textrm{A,e}}(0)P_{\textrm{A,e}}(\tau)}.$
in which one time delay $\tau_{\textrm{A}}$ is set to coincidence. Thus, the second photon B is simultaneously cross-correlated with a photon A and an electron. Now two correlation functions $P_{\textrm{A,B,e}}(\tau)$ and $P_{\textrm{B,e}}(\tau)$ define the shape of the curve, both possibly widened by detector jitters.
For a discrete intensity correlation function, two photons are correlated as a function of the number of heralds between their arrival $q$, comparing the number of successful coincidences between two photons correlated to the same electron ($q=0$) to the number of successful coincidences of two photons correlated to two different electrons ($q\neq 0$) assuming the electrons came at the same time. A successful coincidence is defined by setting a time interval, given by the temporal jitters of the system, in which two particles have to arrive compared to each other.
@@
$g^{(2)}[q]=\frac{N_{\textrm{e,A,B}}[q]N_{\textrm{e}}}{N_{\textrm{e,A}}N_{\textrm{e,B}}}.$ (4)

### v2

@@ Abstract
To test the non-classicality of the radiation process, i.e., whether it is incompatible with classical field theory, we invoke the Cauchy-Schwarz inequality (CSI) [Reid1986] $\gamma=\frac{g_{E\textrm{A}\vee\textrm{B}}^{(2)}(\tau)}{g_{E}^{(2)}g^{(2)}(0)}\leq 1$, assuming a classical radiation process. The inequality (for derivations, see SI) bounds the temporal cross-correlation between the detected electron energy and photon number $g_{E,\textrm{A$\vee$B}}^{(2)}(\tau)$ to the variation of the electron energy loss $g_{E}^{(2)}=\frac{\braket{\Delta E_{el}^{2}}}{\braket{\Delta E_{el}}^{2}}$ and the temporal intensity correlation of photons in a Hanbury-Brown and Twiss setup $g^{(2)}(\tau)$, normalized to $\tau\gg 0$. For electron-photon coincidence, we find a pronounced violation of the CSI ($\gamma>1$, cf. Fig. 4c), which proves the non-classical nature of the process [Clauser1974, Riedinger2016, Meesala2024].
@@
The photon statistics of this radiation are further examined by the temporal intensity correlation $g^{(2)}$ of detectors placed in a Hanbury-Brown Twiss (HBT) setup (see Methods). While classical light shows $g^{(2)}(0)\geq g^{(2)}(\tau)$, this is not the case for non-classical light such as single photons or higher photon number states. The unheralded photon-photon intensity correlation (taken with higher time-resolution, higher-efficiency detectors and a time tagger, see Methods) shows a bunching peak at $\tau=0$ ps [Yanagimoto2025a], inversely proportional to the electron current (cf. Fig. 4d). Following a current-dependent bunching height $g^{(2)}(0)\sim 1+1/(I\tau_{\mathrm{bin}})$, where $I$ is the electron flux rate and $\tau_{\mathrm{bin}}$ is the temporal data bin width, this behavior is expected when single electrons can create multiple photons, either in a direct process, as demonstrated here, or as a result of cascaded excitation of multiple two-level systems leading to incoherent emission [Meuret2015, Fiedler2023, Yuge2023, Sola-Garcia2021a].
@@
The statistics of the non-classical electron-heralded photon state are analyzed using the heralded intensity correlation (see Methods)
@@
in which the detected electrons are energy-filtered to a defined loss peak and used as heralds for the photon state. Thus, the correlation equation relates the occurrence of electron-photon-photon coincidences $N_{\textrm{e,A,B}}$ (cf. Fig. 4a,b) to their single-photon equivalents $N_{\textrm{e,A/B}}$ and the number of heralding electrons $N_{\textrm{e}}$, normalizing $g^{(2)}_{H}(\tau_{\textrm{A}},\tau_{\textrm{B}})$ in this manner to single-photon events ($\tau_{A/B}\approx 0$, $\tau_{B/A}\neq 0$). This uncovers the true threefold coincidences at $\tau_{\textrm{A$\wedge$B}}=0$.
@@
Figure 4e displays $g^{(2)}_{H}(\tau_{\textrm{A}},\tau_{\textrm{B}})$ for single-photon electron-energy losses (m=1{m=1}), as well as the so-called time-averaged heralded intensity correlation $\overline{g}_{H}^{(2)}(\tau)$ reduced to the relative photon delay (inset; see SI for calculation) [Bocquillon2009, Bettelli2010]. The background level and the diagonal feature ($\tau_{\textrm{A$\wedge$B}}\neq 0$) originate from accidental coincidences (cf. Fig. 4a). At $\tau_{\textrm{A$\wedge$B}}\approx 0$, $g^{(2)}_{H}$ drops sharply to near-zero values, as also visible in the inset. Hence, for $m=1$, two photons detected at similar times are generated by two different electrons. Usually, accidental and true heralded photon-photon coincidences cannot be distinguished in the heralded intensity correlation if the coherence time is lower than the respective detector jitters [Bettelli2010, Hockel2011] (see also SI). However, in our case, the electron-heralded photon-photon coincidences remain distinct due to the much better relative photon detection jitters compared to the heralding electron detector (see SI).
A discrete version of the intensity correlation function $g^{(2)}_{H}[q]$ is given by
$g_{H}^{(2)}[q,E_{\textrm{el}}]=\frac{N_{\textrm{e,A,B}}[q,E_{\textrm{el}}]N_{\textrm{e}}(E_{\textrm{el}})}{N_{\textrm{e,A}}(E_{\textrm{el}})N_{\textrm{e,B}}(E_{\textrm{el}})}.$ (4)
The equation substitutes the photon-photon time delay $\tau$ with the number of heralding electrons $q$ by which photon arrivals on detector A are shifted relative to B for cross-correlation. The number of electron-photon coincidences is now counted by only considering a photon detection within a relative time window of $\pm 2.6$ ns (95% coincidence interval in Fig. 2c) to a detected electron (for details, see Methods and SI). The case $q=0$ describes coincidences between two photons heralded by the same electron. By definition, $g_{\textrm{H}}^{(2)}$ approaches unity for large $|q|$. The resulting intensity correlations are shown in Fig. 4f. Without energy filtering (top), a small bunching peak is observed with $g^{(2)}_{H}[0]=1.10$, similar to the unheralded intensity correlation function (cf. Fig. 4d), but with bin sizes corresponding to the electron detector resolution. Filtering to a single-quantum energy exchange ($m=1$), however, leads to a strong anti-bunching behavior with $g^{(2)}_{H}(0)$ suppressed to 0.056. The pronounced anti-bunching dip observed for both the time-dependent and discrete intensity correlation functions confirms the high purity of the heralded single-photon state. Two-photon states should, under ideal conditions, also show prominent antibunching. At present, however, technical aspects, including optical losses, electron spectral mixing, and the statistics of the electron source, limit the ultimate purity of the heralded state (see SI). Such technical limitations can be improved upon by time-gated electron sources and enhanced sample designs.
@@
Quantifying our scheme in terms of electron-heralded photon generation requires a separation between different electron scattering pathways. We observe inelastic electron scattering at the material, as well as the parametric photon generation into broadband spatial optical modes. The unheralded electron spectrum samples a range of coupling constants reaching $g_{0}^{\textrm{EELS}}>0.5$, slightly smaller than the values reported for extended planar structures [Adiv2023]. Notably, in the present experiments, a large fraction of generated excitations populate a single broadband optical mode. With a value of $g_{0}>0.2$, we report the highest coupling of electrons to a detected photonic mode so far. Reaching the strong-coupling regime by establishing unity coupling constant to a single optical mode [Kfir2019, Zhao2025, Xie2025, Karnieli2024, Huang2023] will give direct access to a wide class of new phenomena, such as electron-photon entanglement [Rotunno2023, Rasmussen2024, Henke2025a, Kazakevich2024], photon-induced electron-electron correlations [Kfir2021, Kumar2024a] and the generation of more complex non-Gaussian optical states [Sirotin2024].
@@
In conclusion, we introduce electron-heralded quantum states of light, identifying hallmark features such as antibunching photon statistics and multiphoton-electron correlations. The scheme will facilitate future electron-heralded high-order Fock states, based on further technical optimizations and strong single-mode coupling. Combined with photon or electron state operations, the generation of a wide variety of complex quantum states comes into reach. Broadening the scope of free-electron quantum optics, our results experimentally establish electron beams as a tailored quantum optical resource.
@@ Methods
The photonic chip with a $\text{Si}_{3}\text{N}_{4}$ waveguide embedded in $\text{SiO}_{2}$ is fabricated via the Damascene process [Liu2021] onto a Si substrate. The waveguide (width x height $2.2\mu\text{m}$ x $780$ nm) has a separation between the bus waveguides, which connect the interaction region with the chip-to-fiber couplers, of around $125\,\mu\text{m}$. The longest path at which electrons can intersect with the waveguide is around $80\penalty 10000\ \mu\text{m}$ long. However, the curvature of the waveguide, as well as slight angular deviations between the waveguide and the electron beam, reduce the effective interaction length to around $40\penalty 10000\ \mu\text{m}$. The waveguide has no top cladding, allowing for the optical modes to reach into the vacuum and interact with the electron beam. An ITO (indium tin oxide) layer of <20<20 nm thickness (index of refraction $n=1.99$ at 1550 nm measured at ITO films deposited under similar conditions) on top of the waveguide reduces charging in the presence of an electron beam. The chip is mounted on a custom-made TEM holder [Henke2021] such that the waveguide sits in parallel to the electron beam. The optical outputs of the symmetrically-built chip are connected to two short pieces of ultra-high numerical aperture (UHNA-7) fiber, which are spliced to single-mode fibers (SMF-28). The single-mode fibers are sent through the hollowed-out TEM holder and a custom vacuum fiber feedthrough. As the photon generation is directional, the optical detection setup is spliced to the SMF-28 fiber, which is connected to the lower output port.
@@
The experiments are conducted at the Göttingen UTEM (JEOL JEM 2100F) with a Schottky field-emission electron source operated in a continuous (thermal-field) mode. The resulting electron beam shows an energy spread of around 0.6 eV at an electron energy of 100 keV. The current is reduced to around 1.4 pA by adjusting the emitter temperature and by using a condenser aperture of $40\penalty 10000\ \mu\text{m}$ diameter to not oversaturate the electron detector. The experiments are conducted in low-magnification STEM (scanning transmission electron microscopy) mode, enabling the positioning and scanning of the focused electron beam (30 nm beam diameter, 1.1 mrad half convergence angle) in the sample plane. The scanning resolution is on the order of 30 nm. All measurements described in the main text are conducted with a stationary electron beam, which is checked and manually readjusted for an optimal combination of photon generation rate and electron transmission onto the detector every 10-30 min. Transmission losses on the electron side arise from charging effects along the waveguide surface, distorting and shifting the electron beam, which reduces incoupling into the 5 mm spectrometer entrance aperture. Post-measurement calculations show an electron transmission efficiency of around 60-70%. For the scanned data, the electron beam is scanned multiple times across the same region (width x height 200x500nm, 300 ms integration per pixel, pixel size $\sim 40$ nm). The scanned data is used to measure the position-dependent change of the coupling constant.
@@
The generated photons are sent through the waveguide and optical fibers to the outside of the TEM holder. The fiber end is spliced to a 50/50 beam splitter (Thorlabs TW1550R5A1), which distributes the photons to two cooled avalanche photodiodes (ID Quantique ID230). As the generated light is measured to be around $1300-1400\text{nm}$, we assume slightly higher losses and a deviation from the 50/50 coupling ratio at the beam splitter. The count rates indicate a splitting ratio of 52/48. The avalanche photodiodes are set to $25\%$ detection efficiency and $20\mu\text{s}$ dead time, which leads to 250 and 300 photons/s intrinsic dark counts, respectively, for diode A and B. An evaluation of the transmission and detection efficiency shows an approximate efficiency of detecting a generated photon of 2% per diode. The optical spectrum is measured using a 10 nm flat-top bandpass filter (WL Photonics, tunability from 1200 to 1600nm), which is set between the TEM holder and the beam splitter, connected to the optical fibers via FC/APC connectors. The transmission band was moved incrementally between 1270 and 1570nm and the count rate on the detector was measured for every position (see SI). The spectrum shows a central wavelength of 1380 nm and a bandwidth of 100 nm (cf. Fig3e). The unheralded photon-photon intensity correlation is measured with a combination of superconducting nanowire single-photon detectors (Single Quantum) and a time tagger (Swabian Instruments, Time Tagger Ultra). The nanowire detectors show a detection efficiency $80\%$ and a time resolution <20<20 ps and, therefore, allow for sharper temporal correlation when operated in combination with a Tagger unit with an RMS jitter of 9 ps.
@@
The electrons are separated in energy via a spectrometer (CEOS GmbH, CEFID) and detected on a hybrid pixel detector based on 4 Timepix chips with 256x256 readout pixels each (Amsterdam Scientific Instruments, EM CheeTah T3). The event-based data stream consists of electron hits (0.03-eV energy bin per pixel on the dispersion axis, 1.56-ns timing bin) and synchronized signals from two time-to-digital converters (260-ps timing bin) at which the photon detector signals arrive. Both the electron and photon time measurements are synchronized to the same global readout clock. This allows for the temporal correlation of the electron and photon arrival. Further post-acquisition steps are conducted to improve temporal and energy resolution on the electron side: First, a calibration of the individual pixel response times is done with regard to a mean detector response to enhance time resolution [Feist2022]. Second, single-particle clustering of the multiple activated pixels/hits (3.4 hits per cluster at 100-keV beam energy, maximum cluster size of 10) gives access to individual electron events. Their localization follows the arithmetic mean position and the earliest arrival time inside a single cluster. A comparison between external current measurements and event counting on the detector shows an average of 3.4 excited pixels per electron. Last, time-dependent jitters of the electron beam position are reduced on a 10-s interval by correcting the position of the zero-loss peak maximum.
@@
For the temporal correlation between electrons and photons, every detected electron is tagged with the relative time delay $\tau_{\textrm{A/B}}=t_{\textrm{A/B}}-t_{\textrm{el}}$ to the temporally closest photons on detector A and detector B (absolute propagation timing offsets are subtracted). Hence, the resulting electron-photon-photon correlation data contains the electron kinetic energy $E$ (calibrated by the pixel position) and the (relative) arrival times of all three particles. For scanned data, the electron beam position is also included.
A thorough analysis and quantitative fit of event rates for this three-dimensional event histogram ($E,\tau_{\textrm{A}},\tau_{\textrm{B}}$) can be found in the SI. For easier access, individual cuts along or integrals over a specific axis are also analyzed and fitted independently. Generally, this procedure may include heralded multi-indexing of photons within the relevant time interval of 100 ns. For the calculation of the discrete heralded intensity correlation $g_{H}^{(2)}[q,E_{\textrm{el}}]$, as well as Fig. 2d, photons were assigned uniquely, only to the temporarily closest electron.
@@
All data besides the uncorrelated spectrum (Fig. 2) is analysed after filtering the electrons to events with a maximum electron-photon time delay of 100 ns, which strongly reduces computation time.
@@
The statistics of photons and the proof for single photon states are often given via the intensity correlation measured in a Hanbury-Brown and Twiss setup in which the signal from two single-photon detectors is cross-correlated and normalized to large time delays [Signorini2020, Migdall2013]
@@
Here, $P_{\textrm{A,B}}(\tau)$ is the probability of detector A and B being excited at a temporal delay $\tau$ to each other, and $P_{\textrm{i}}$, i=A,B is the probability of detector excitation A or B.
As the electrons do not arrive uniformly, their statistical arrival distribution plays a role in the unheralded photon-photon intensity correlation. To remove the electron influence, an electron-heralded intensity correlation is employed. Heralded intensity correlation follows [Bocquillon2009]
@@
in which the photon-photon cross-correlation $P_{\textrm{A,B,e}}(t_{\textrm{A}},t_{\textrm{B}},t_{\textrm{e}})$ is analysed in combination with the detection of a heralding particle $e$ (see SI for more information). While the numerator $P_{\textrm{A,B,e}}$ becomes a function of all three particle arrival times, the denominators convert into correlation functions between one photon and the heralding electron $P_{\textrm{A/B,e}}$ as a function of electron-photon time delays $\tau_{\textrm{A/B}}=t_{\textrm{A/B}}-t_{\textrm{e}}$. The function is again normalized to large photon-photon time delays, but under the condition that either photon A or photon B is in coincidence with a herald $e$. For electron-heralding intensity correlation, the equation is rewritten in terms of correlation counts $N$ instead of coincidence probabilities, giving rise to Equation 3.
Due to the addition of a third particle, the heralded intensity correlation is a function of two time delays $\tau_{\textrm{A}}$ and $\tau_{\textrm{B}}$. A reduction to one single variable is done in time-averaged intensity correlation [Bocquillon2009, Bettelli2010]
$\overline{g}^{(2)}(\tau_{\textrm{A}}=0,\tau=\tau_{\textrm{B}})=\frac{P_{\textrm{A,B,e}}(\tau)P_{\textrm{e}}}{P_{\textrm{A,e}}(0)P_{\textrm{A,e}}(\tau)}.$
in which one time delay $\tau_{\textrm{A}}$ is set to coincidence. Thus, the second photon B is simultaneously cross-correlated with a photon A and an electron.
For a discrete intensity correlation function, two photons are correlated as a function of the number of heralding electrons $q$ by which a photon on detector A is shifted relative to a photon on detector B during cross-correlation. The equation, therefore, compares the number of successful coincidences between two photons correlated to the same electron ($q=0$) to the number of successful coincidences of two photons correlated to two different electrons ($q\neq 0$), assuming the electrons came at the same time. A successful coincidence is defined by setting a time interval of 2.9 ns, given by the temporal jitters of the system, in which two particles have to arrive compared to each other.
@@
$g^{(2)}[q,E_{\textrm{el}}]=\frac{N_{\textrm{e,A,B}}[q,E_{\textrm{el}}]N_{\textrm{e}}(E_{\textrm{el}})}{N_{\textrm{e,A}}(E_{\textrm{el}})N_{\textrm{e,B}}(E_{\textrm{el}})}.$ (5)
@@
Generally, all (heralded) intensity correlation functions require noise removal in the individual correlation functions to be normalized for large time delays. However, high signal-to-noise ratios in detector counts, as experienced in this experiment, eliminate the necessity for noise removal. All shown intensity correlation measurements were conducted without background removal. Furthermore, the heralding electrons are generally filtered to a defined energy loss to achieve photon number state analysis on the heralded photon side. The requirement of a fixed electron energy loss value is clarified by adding Eel as a further variable in the intensity correlation functions.
Data availability statement
The code and data used to produce the figures in this work will be published in the repository Zenodo upon publication of the paper.